OpenStudy (anonymous):

How hard do you need to throw to get the projectile to the desired location at three different angles; 30,45,60. x1=10m y1=20m

OpenStudy (anonymous):

I believe, you have read those kinematics equations in projectile motion.

OpenStudy (anonymous):

Did you try writing the equations along x , along y and solving them ? I think, it'll be solved that way. We know x, y, g and theta . We need to find 'u' and we don't know 't'. 2 equations with 2 variables, it'll be solved ..

OpenStudy (anonymous):

What do you mean when you say writing the equations along x and along y? Do you mean x=x0+vxo+1/2axt^2 and y=y0+Vyo+1/2ayt^2 ?

OpenStudy (anonymous):

ya, in this case ax = 0. So, you can easily get t expression fro there and put that in eq for y.

OpenStudy (anonymous):

ok and I would use the y (20m) for the height (y)?

OpenStudy (anonymous):

yes, as given in the q, x1=10, and y1=20

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

:) Do convey problems if you face any by chance.

OpenStudy (anonymous):

Will do. Thank you

OpenStudy (anonymous):

oh, you just thanked me twice. You're welcome !